# Stoichiometry Representation Learning with Polymorphic Crystal Structures

### v1

@@ 4.2 Probabilistic Stoichiometry Encoder
Deterministic Representation. After obtaining the structural representation $\mathbf{z}^{b}$, we also compute the stoichiometry representation from the stoichiometry graph $\mathcal{G}^{a}$ as follows:
$\mathbf{z}^{a}=\text{Pooling}(\mathbf{Z}^{a}),\,\,\,\,\mathbf{Z}^{a}=f^{a}(\mathbf{X}^{a},\mathbf{A}^{a}),$ (2)
@@
where $\mathbf{Z}^{a}\in\mathbb{R}^{n_{e}\times F}$ is a matrix whose each row indicates the representation of each element in a stoichiometry, $\mathbf{z}^{a}\in\mathbb{R}^{F}$ indicates the stoichiometry representation of a crystal, and $f^{a}$ is a GNN-based stoichiometry encoder. By utilizing GNNs, the stoichiometry encoder effectively learns intricate relationships and chemical environments related to elements, thereby enhancing the stoichiometry representation in a systematic manner Goodall & Lee (2020). For the stoichiometry encoder $f^{a}$, we adopt GCNs Kipf & Welling (2016) with jumping knowledge Xu et al. (2018), and weighted sum pooling with the compositional ratio (i.e., $\mathcal{R}$ in Section 3.1) is used as the pooling function.
@@
One straightforward approach for injecting structural information into the stoichiometry representation would be adopting the idea of recent multi-modal contrastive learning approaches, which have been widely known to maximize the mutual information between heterogeneous modality inputs (two modalities in our case: stoichiometry and structure) Radford et al. (2021); Stärk et al. (2022). However, such a naive adoption fails to capture the polymorphic nature of crystallography: A single stoichiometry can result in multiple distinct structures due to the diverse atomic arrangements, leading to significantly different physical, and chemical properties Bernstein (2020). That is, the relationship between the representations $\mathbf{z}^{a}$ and $\mathbf{z}^{b}$ constitutes a one-to-many mapping rather than a one-to-one mapping, leading to inherent uncertainties in the stoichiometry representation $\mathbf{z}^{a}$.
@@
Probabilistic Representation. To this end, we propose to learn a probabilistic representation of stoichiometry $\mathbf{z}^{a}$, which naturally exhibits uncertainties of the representation, inspired by the recent Hedge Instance Embeddings (HIB) Oh et al. (2018). The main idea here is to learn the Gaussian representation of stoichiometry, which reveals the distribution of polymorphic structures $\mathcal{P}^{a}$ in representation space. Intuitively, the variance of this distribution reflects the range of diversity within these structures, giving us an idea of how well the representation adheres to the principle of preservativity. More formally, we model each stoichiometry as a parameterized Gaussian distribution with learnable mean vectors and diagonal covariance matrices as follows:
$p(\tilde{\mathbf{z}}^{a}|\mathbf{X}^{a},\mathbf{A}^{a})\sim\mathcal{N}(\mathbf{z}_{\mu}^{a},\mathbf{z}_{\sigma}^{a}),\,\,\,\,\text{where}\,\,\,\,\mathbf{z}_{\mu}^{a}=f^{a}_{\mu}(\mathbf{Z}^{a}),\,\,\,\,\mathbf{z}_{\sigma}^{a}=f^{a}_{\sigma}(\mathbf{Z}^{a}).$ (3)
@@
Here, $\mathbf{z}_{\mu}^{a},\mathbf{z}_{\sigma}^{a}\in\mathbb{R}^{F}$ denote the mean vector and the diagonal entries of the covariance matrix, respectively, and $f^{a}_{\mu}$ and $f^{a}_{\sigma}$ refer to the modules responsible for calculating the mean and diagonal covariance matrices, respectively. During training, we adopt the re-parameterization trick Kingma & Welling (2013) to obtain samples from the distribution, i.e., $\tilde{\mathbf{z}}^{a}=diag(\sqrt{\mathbf{z}_{\sigma}^{a}})\cdot\epsilon+\mathbf{z}_{\mu}^{a}$, where $\epsilon\sim\mathcal{N}(0,1)$. While mean and variance are obtained from the shared $\mathbf{Z}^{a}$, we utilize different attention-based set2set pooling functions for $f^{a}_{\mu}$ and $f^{a}_{\sigma}$ Vinyals et al. (2015), since the attentive aspects involved in calculating the mean and variance should be independent from each other. We provide further details on the probabilistic stoichiometry encoder in Appendix B.2.
4.3 Model Training via Representation Alignment
@@
To incorporate the structural information into the stoichiometry representation, we define a matching probability between the stoichiometry graph $\mathcal{G}^{a}$ and its corresponding set of polymorphic crystal structural graphs $\mathcal{P}^{\mathcal{G}^{a}}$ in the Euclidean space as follows:
@@
$p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a}})\approx\sum_{p\in\mathcal{P}^{a}}\frac{1}{J}\sum^{J}_{j=1}{\text{sigmoid}\big{(}-c{\|\tilde{\mathbf{z}}_{j}^{a}-\mathbf{z}_{p}^{b}\|}_{2}+d\big{)}},$ (4)
@@
where $\tilde{\mathbf{z}}_{j}^{a}$ is the sampled stoichiometry representation (Section 4.2), $\mathbf{z}_{p}^{b}$ is the structural graph representation (Section 4.1), $c,d>0$ are parameters learned by the model for soft threshold in the Euclidean space, $J$ is the number of samples sampled from the distribution, and $\text{sigmoid}(\cdot)$ is the sigmoid function. Moreover, $m$ is the indicator function of value 1 if $\mathcal{P}^{\mathcal{G}^{a}}$ is the set of polymorphic structures corresponding to $\mathcal{G}^{a}$ and 0 otherwise. Then, we apply the soft contrastive loss Oh et al. (2018); Chun et al. (2021) as:
$\mathcal{L}_{\text{con}}=\begin{cases*}-\log{p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a\prime}})},&if $a=a^{\prime}$,\\ -\log{(1-p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a\prime}}))},&otherwise.\\ \end{cases*}$ (5)
@@
Intuitively, the above loss aims to minimize the distance between a sampled stoichiometry representation and its associated polymorphic structural representations, while maximizing the distance between others. By doing so, PolySRL learns a probabilistic stoichiometry representation that considers the structural information and its associated uncertainties, which tend to increase when multiple structures are associated with a single stoichiometry, i.e., polymorphism.
@@
In addition to the soft contrastive loss, we utilize a KL divergence loss between the learned stoichiometry distributions and the standard normal distribution $\mathcal{N}(0,1)$, i.e., $\mathcal{L}_{\text{KL}}=\text{KL}(p(\tilde{\mathbf{z}}^{a}|\mathbf{X}^{a},\mathbf{A}^{a})~{}\|~{}\mathcal{N}(0,1))$, which prevents the learned variances from collapsing to zero. Therefore, our final loss for model training is given as follows:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{con}}+\beta\cdot\mathcal{L}_{\text{KL}},$ (6)
@@
where $\beta$ is the hyperparameter for controlling the weight of the KL divergence loss. During the inference, we use the mean vector $\mathbf{z}_{\mu}^{a}$ as the stoichiometry representation and the geometric mean of diagonal covariance matrices $\mathbf{z}_{\sigma}^{a}$ as uncertainty Chun et al. (2021).
5 Experiments
5.1 Experimental Setup
Datasets. For training PolySRL, we collect 80,162 unique stoichiometries and their corresponding 112,183 DFT-calculated crystal structures from Materials Project (MP) website  [^2] https://materialsproject.org/. However, since DFT-calculated properties often deviate from real-world wet-lab experimental properties  (Jha et al., 2019), we primarily evaluate PolySRL using wet-lab experimental datasets. Specifically, we use publicly available datasets containing experimental properties of stoichiometry, including Band Gap Zhuo et al. (2018), Formation Enthalpies Kim et al. (2017), Metallic Morgan (2018), and ESTM Na & Chang (2022). Moreover, we conduct experiments on seven Matbench Dunn et al. (2020) datasets that are related to DFT-calculated properties. We provide further details on the datasets in Appendix C.
Baseline Methods. Since PolySRL is the first work that learns stoichiometry representation without any label information, we construct competitive baseline models from other domains. Rand init. refers to a randomly initialized stoichiometry encoder without any training process. GraphCL You et al. (2020) learns the stoichiometry representation based on random augmentations on the stoichiometry graph $\mathcal{G}^{a}$, without utilizing structural information. MP Band G. and MP Form. E. learn the stoichiometry representation by predicting the DFT-calculated properties, which are available in MP database 2, i.e., band gap and formation energy per atom, respectively. 3D Infomax Stärk et al. (2022) learns stoichiometry representation by maximizing the mutual information between stochiometry graph $\mathcal{G}^{a}$ and structural graph $\mathcal{G}^{b}$ with NTXent (Normalized Temperature-scaled Cross Entropy) loss Chen et al. (2020). We provide further details on baseline methods in Appendix D. In addition, we also compare PolySRL with supervised stoichiometry representation learning methods, i.e., Roost Goodall & Lee (2020) and CrabNet Wang et al. (2021) in Appendix F.2.
Evaluation Protocol. We first train all models in an unsupervised manner without any use of wet-lab experimental data. Then, we evaluate PolySRL in two evaluation schemes, i.e., representation learning and transfer learning. We further provide the detailed evaluation protocols in Appendix E.
@@ 5.2 Empirical Results
Representation Learning. In Table 1, we have the following observations: 1) Comparing the baseline methods that take into account structural information (Str. ) with those that do not (Str. ), we find out that utilizing structural information generally learns more high-quality stoichiometry representations. This is consistent with the established knowledge in crystallography, which emphasizes that structural details, including crystal structure and symmetry, play a crucial role in determining a wide range of physical, chemical, and mechanical properties Bernstein (2020); Braga et al. (2009). 2) Moreover, we observe PolySRL outperforms baseline methods that overlook polymorphism in their model design. This highlights the significance of our probabilistic approach, which not only offers insights into polymorphism-related uncertainties but also yields high-quality representations. 3) On the other hand, we notice that utilizing DFT-calculated values contributes to the model’s understanding of a specific target property (see Prop. ). For instance, when the model is trained with a DFT-calculated band gap (i.e., MP Band G.), it surpasses all other models when predicting experimental band gap values. This highlights that knowledge acquired from DFT-calculated properties can be applied to wet-lab experimental datasets. However, these representations are highly tailored to a particular target property, which restricts their generalizability for diverse tasks. We also provide empirical results on Matbench datasets that contain DFT-calculated properties in Appendix F.1 and transfer learning scenarios in Appendix F.2.

### v2

@@ 2.3 Probabilistic Composition Encoder $f^{a}$
Deterministic Representation. After obtaining the structural representation $\mathbf{z}^{b}$, we also compute the compositional representation from the composition graph $\mathcal{G}^{a}$ as follows:
$\small\mathbf{z}^{a}=\text{Pooling}(\mathbf{Z}^{a}),\,\,\,\,\mathbf{Z}^{a}=f^{a}(\mathbf{X}^{a},\mathbf{A}^{a}),$ (2)
@@
where $\mathbf{Z}^{a}\in\mathbb{R}^{n_{e}\times F}$ is a matrix whose each row indicates the representation of each element in a composition, $\mathbf{z}^{a}\in\mathbb{R}^{F}$ indicates the compositional representation of a crystal, and $f^{a}$ is a GNN-based composition encoder. By utilizing GNNs, the composition encoder effectively learns intricate relationships and chemical environments related to elements, thereby enhancing the compositional representation in a systematic manner [9]. For the composition encoder $f^{a}$, we adopt GCNs [26] with jumping knowledge [27], and weighted sum pooling with the compositional ratio (i.e., $\mathcal{R}$ in Section 2.1.1) is used as the pooling function.
@@
One straightforward approach for injecting structural information into the compositional representation would be adopting the idea of recent multi-modal contrastive learning approaches, which have been widely known to maximize the mutual information between heterogeneous modality inputs (two modalities in our case: composition and structure) [14, 15]. However, such a naive adoption fails to capture the polymorphic nature of crystallography: A single composition can result in multiple distinct structures due to the diverse atomic arrangements, leading to significantly different physical, and chemical properties [16]. That is, the relationship between the representations $\mathbf{z}^{a}$ and $\mathbf{z}^{b}$ constitutes a one-to-many mapping rather than a one-to-one mapping, leading to inherent uncertainties in the compositional representation $\mathbf{z}^{a}$.
@@
Probabilistic Representation. To this end, we propose to learn a probabilistic representation of composition $\mathbf{z}^{a}$, which naturally exhibits uncertainties of the representation, inspired by the recent Hedge Instance Embeddings (HIB) [21]. The main idea here is to learn the Gaussian representation of composition, which reveals the distribution of polymorphic structures $\mathcal{P}^{a}$ in representation space. Intuitively, the variance of this distribution reflects the range of diversity within these structures, thereby providing an insight into the level of uncertainty associated with composition. More formally, we model each composition as a parameterized Gaussian distribution with learnable mean vectors and diagonal covariance matrices as follows:
$\small p(\tilde{\mathbf{z}}^{a}|\mathbf{X}^{a},\mathbf{A}^{a})\sim\mathcal{N}(\mathbf{z}_{\mu}^{a},\mathbf{z}_{\sigma}^{a}),$ (3)
@@
where $\mathbf{z}_{\mu}^{a}=f^{a}_{\mu}(\mathbf{Z}^{a})$ and $\mathbf{z}_{\sigma}^{a}=f^{a}_{\sigma}(\mathbf{Z}^{a})$. Here, $\mathbf{z}_{\mu}^{a},\mathbf{z}_{\sigma}^{a}\in\mathbb{R}^{F}$ denote the mean vector and the diagonal entries of the covariance matrix, respectively, and $f^{a}_{\mu}$ and $f^{a}_{\sigma}$ refer to the modules responsible for calculating the mean and diagonal covariance matrices, respectively. During training, we adopt the re-parameterization trick [28] to obtain samples from the distribution, i.e., $\tilde{\mathbf{z}}^{a}=diag(\sqrt{\mathbf{z}_{\sigma}^{a}})\cdot\epsilon+\mathbf{z}_{\mu}^{a}$, where $\epsilon\sim\mathcal{N}(0,1)$. While mean and variance are obtained from the shared $\mathbf{Z}^{a}$, we utilize different attention-based set2set pooling functions for $f^{a}_{\mu}$ and $f^{a}_{\sigma}$ [29], since the attentive aspects involved in calculating the mean and variance should be independent of each other. We provide further details on the probabilistic composition encoder in Appendix B.2.
2.4 Model Training via Sampled Representation Alignment
@@
To incorporate the structural information into the compositional representation, we define a matching probability between the composition graph $\mathcal{G}^{a}$ and its corresponding set of polymorphic crystal structural graphs $\mathcal{P}^{\mathcal{G}^{a}}$ in the Euclidean space as follows:
@@
$\small p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a}})\approx\sum_{p\in\mathcal{P}^{\mathcal{G}^{a}}}\frac{1}{J}\sum^{J}_{j=1}{\text{sigmoid}\big{(}-c{\|\tilde{\mathbf{z}}_{j}^{a}-\mathbf{z}_{p}^{b}\|}_{2}+d\big{)}},$ (4)
@@
where $\tilde{\mathbf{z}}_{j}^{a}$ is the sampled compositional representation, $\mathbf{z}_{p}^{b}$ is the structural graph representation, $c,d>0$ are parameters learned by the model for soft threshold in the Euclidean space, $J$ is the number of samples sampled from the distribution, and $\text{sigmoid}(\cdot)$ is the sigmoid function. Moreover, $m$ is the indicator of value 1 if $\mathcal{P}^{\mathcal{G}^{a}}$ is the set of polymorphic structures corresponding to $\mathcal{G}^{a}$ and 0 otherwise.
Then, we apply the soft contrastive loss [21, 22] as:
$\small\mathcal{L}_{\text{con}}=\begin{cases*}-\log{p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a\prime}}}),&if $a=a^{\prime}$,\\ -\log{(1-p(m|\mathcal{G}^{a},\mathcal{P}^{\mathcal{G}^{a\prime}}}),&otherwise.\\ \end{cases*}$ (5)
@@
Intuitively, the above loss aims to minimize the distance between a sampled compositional representation and its associated polymorphic structural representations, while maximizing the distance between others. By doing so, PCRL learns a probabilistic compositional representation that considers the structural information and its associated uncertainties, which tend to increase when multiple structures are associated with a single composition, i.e., polymorphism.
@@
In addition to the soft contrastive loss, we utilize a KL divergence loss between the learned composition distributions and the standard normal distribution $\mathcal{N}(0,1)$, i.e., $\mathcal{L}_{\text{KL}}=\text{KL}(p(\tilde{\mathbf{z}}^{a}|\mathbf{X}^{a},\mathbf{A}^{a})\leavevmode\nobreak\ \|\leavevmode\nobreak\ \mathcal{N}(0,1))$, which prevents the learned variances from collapsing to zero. Therefore, our final loss for model training is given as follows:
$\small\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{con}}+\beta\cdot\mathcal{L}_{\text{KL}},\vspace{-0.5ex}$ (6)
@@
where $\beta$ is the hyperparameter for controlling the weight of the KL divergence loss. During the inference, we use the mean vector $\mathbf{z}_{\mu}^{a}$ as the compositional representation and the geometric mean of diagonal covariance matrices $\mathbf{z}_{\sigma}^{a}$ as uncertainty [22].
2.5 Datasets
  • Materials Project [23] is an openly accessible database that provides material properties calculated using density functional theory (DFT). We have gathered 80,162 distinct compositions along with their corresponding 112,183 crystal structures computed using DFT, with up to 32,021 compositions having multiple potential structures.
  • Band Gap [30] dataset comprises experimentally determined band gap properties for non-metallic materials. It encompasses 2,482 distinct compositions and a total of 3,895 experimental band gap values. Within this dataset, 1,413 instances of duplicate experimental band gap measurements for compositions were identified. Consequently, our task involves predicting the band gap properties for these 2,482 compositions, with the average value being computed in cases where duplicate experimental results exist for a given composition.
  • Formation Enthalpies [31] dataset consists of experimentally determined formation enthalpy values for intermetallic phases and other inorganic compounds. It includes 1,141 unique compositions and a total of 1,276 experimental formation enthalpy values. Within this dataset, 135 cases of duplicate experimental formation enthalpy measurements for compositions were identified. Therefore, our objective is to predict the formation enthalpy properties for these 1,141 compositions, calculating the average value when duplicate experimental results are present for a particular composition. We report MAE values multiplied by a factor of 10 for clear interpretation during all experiments.
  • Metallic [32] dataset contains reduced glass transition temperature (Trg) for 584 unique metallic alloys. We report MAE values multiplied by a factor of 10 for clear interpretation during all experiments.
  • ESTM 300 K [33] dataset contains various properties of 368 thermoelectric materials that are measured in the temperature range of 295 K to 305 K, which is widely recognized as room temperature in chemistry. Among the properties, we mainly target electrical conductivity ($S/m$), thermal conductivity ($W/mK$), and Seebeck coefficient ($\mu V/K$). Regarding electrical conductivity and thermal conductivity, we apply a logarithmic scaling to the target values because they exhibit significant skewness. Additionally, for the Seebeck coefficient, we use min-max scaling on the target values due to their wide range and report MAE values multiplied by a factor of 10 for clear interpretation during all experiments. When calculating the figure of merit ($Z\bar{T}$) with predicted properties, we reverse the scaling to return the original scale and then compute it.
  • ESTM 600 K [33] dataset contains various properties of 188 thermoelectric materials that are measured in the temperature range of 593 K to 608 K, which is widely recognized as high temperature in chemistry. The properties we are targeting and the preprocessing steps applied are identical to those used for the ESTM 300 K dataset.
In addition to the wet-lab experimental datasets, we use the following seven Matbench datasets that contain properties from DFT calculation.
  • Castelli Perovskites [34] dataset contains formation energy of Perovskite cell of 18,928 materials.
  • Refractive Index [23] dataset contains a refractive index of 4,764 materials, provided in MP database.
  • Shear Modulus [23] dataset contains shear modulus of 10,987 materials, provided in MP database.
  • Bulk Modulus [23] dataset contains bulk modulus of 10,987 materials, provided in MP database.
  • Exfoliation Energy [35] dataset contains exfoliation energy 636 materials.
  • MP Band gap [23] dataset contains band gap of 106,113 materials, provided in MP database.
  • MP Formation Energy [23] dataset contains formation energy per atom in 132,752 materials, provided in MP database.
Following previous work [24, 9], we choose the target value associated with the lowest formation enthalpy for duplicate compositions found in both the MP datasets, while we use the mean of the target values for other datasets.
2.6 Baseline Methods
  • Rand init. refers to the randomly initialized composition encoder without any training process.
  • GraphCL [36] is a general graph-level contrastive learning strategy that uses random augmentation to construct positive and negative samples. In this paper, it learns the compositional representation based on the random augmentation on the composition graph $\mathcal{G}^{a}$, without utilizing structural information. For the $n$-th data in the minibatch ($N$ data points), the loss function is defined as follows follows:
  • MP Band G. and MP Form. E. learn the compositional representation by predicting the DFT-calculated properties, i.e., band gap and formation energy per atom, respectively. More formally, model is trained with MAE loss for $n$-th data point in the minibatch ($N$ data points) as follows:
    $l_{n}=|Y_{n}-\hat{Y}_{n}|,\$ (8)
  • 3D Infomax [15] proposes to enhance model prediction on 3D molecular graphs by integrating 3D information of the molecules in its latent representations. Instead of 2D molecular graphs, we learn the representation of composition graph $\mathcal{G}^{a}$ by maximizing the mutual information with structural graph $\mathcal{G}^{b}$. More specifically, we train the model with NTXent (Normalized Temperature-scaled Cross Entropy) loss [37], which is defined for $n$-th data point in minibatch of size $N$ as follows:
Even though the primary focus of this paper is to introduce training strategies for composition encoders without any label information, we also conduct a comparative analysis of our proposed approach with previous supervised compositional representation learning methods [9, 24]. Note that these works propose sophisticated model architectures for compositional representation learning, not training strategy.
  • Roost [9] first proposes to utilize GNNs for compositional representation learning by presenting composition as a fully connected graph, whose nodes are unique elements in composition. This approach allows the model to acquire distinct and material-specific representations for each element, enabling it to capture physically meaningful properties and interactions.
  • CrabNet [24] designs a Transformer self-attention mechanism [38] to adaptively learn the representation of individual elements based on their chemical environment.
2.7 Evaluation Protocol
Evaluation Metrics. We mainly compare the methods in terms of Mean Absolute Error (MAE) following previous work [9]. Moreover, we provide the model performance in terms of $R^{2}$ in Appendix C, which provides an intuitive measure of the fraction of the overall variance in the data that the model can account for.
Representation Learning. For representation learning scenarios, we fix the model parameters (i.e., $f^{a}$, $f_{\mu}^{a}$, and $f_{\sigma}^{a}$) and train a three-layer MLP head with LeakyReLU non-linearity to evaluate the composition obtained by various models. Following previous works [39, 36], we train the MLP head with Adam optimizer with a fixed learning rate of 0.001 for 300 epochs.
3 Results
3.1 Quantitative Results
@@
Representation Learning. In Table 1, we have the following observations: 1) We find out that utilizing structural information (Str. ) generally learns more high-quality compositional representations compared with those that overlook the structural information (Str. ). This is consistent with the established knowledge in crystallography, which emphasizes that structural details, including crystal structure and symmetry, play a crucial role in determining a wide range of physical, chemical, and mechanical properties [16, 40]. 2) Moreover, we observe PCRL outperforms baseline methods that overlook polymorphism in their model design (Poly. ). This highlights the significance of our probabilistic approach, which not only offers insights into polymorphism-related uncertainties but also yields high-quality representations. 3) On the other hand, we notice that utilizing DFT-calculated values contributes to the model’s understanding of a specific target property (see Prop. ). For instance, when the model is trained with a DFT-calculated band gap (i.e., MP Band G.), it surpasses all other models when predicting specifically the experimental band gap values. This highlights that knowledge acquired from DFT-calculated properties can be helpful in predicting the properties obtained from wet-lab experiments. However, these representations are highly tailored to a particular target property, which restricts their generalizability for diverse tasks. We provide visualization of representation and statistical significance of improvement in Appendix C.1 and C.2, respectively.